OpenStudy (anonymous):

the circle (x+1)^2+(y-4)^2=26 intersects y axis at?

OpenStudy (anonymous):

The graph will intersect the y-axis when x=0. Substituting x=0 into the equation of the circle leaves you with \((0+1)^2+(y-4)^2=26 \implies 1+(y-4)^2=26 \implies (y-4)^2=25\). Can you take things from here and solve for y? :-)

OpenStudy (anonymous):

ok will try

OpenStudy (anonymous):

1 is (0.9) whats the other @ChristopherToni

OpenStudy (anonymous):

can u answe r @ChristopherToni

myininaya (myininaya):

\[(y-4)^2=25 \\ \text{ gives two equations }\]

myininaya (myininaya):

\[y-4= \pm \sqrt{25}\]

OpenStudy (anonymous):

then

OpenStudy (anonymous):

thanks i got that

myininaya (myininaya):

you already solved one of them so solve the one you didn't solve yet

OpenStudy (anonymous):

ty:)

myininaya (myininaya):

np

OpenStudy (anonymous):

the answer is (9,0) and (-1,0) right?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

y-4=5 or y-4=-5 y=9 or y=-1 well almost you let x=0 so the points are (0,9) and (0,-1)

OpenStudy (anonymous):

yeh sorry its (0,9) and (0,-1)

myininaya (myininaya):

anyways good job :)

OpenStudy (anonymous):

ty:)